ok i'm stuck
(a-1)^1/3=-3

Answer by stanbon(75887)

(a-1)^1/3=-3
---------------------
Raise both sides to the 3rd power to get:
(a-1) = (-3)^3
a-1 = -27
a = -26
=============
Cheers,
Stan H.